# Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories

### v2

@@ Abstract
The existence of Quantum Many-Body Scars has been established across different quantum many-body systems, including spin-1/2 Quantum Link Models. We systematically identify scars for pure gauge theories with arbitrarily large integer spin $S$ in $2+1$D, concretely for Truncated Link Models, where the electric field is truncated rather than replaced by spins. Through an explicit analytic construction, we show that the presence of scars is widespread. We confirm these findings numerically for small truncated spin and $S=1$ Quantum Link Models. The proposed analytic construction establishes the presence of scars far beyond volumes and spins that can be probed with existing numerical methods.
@@
Introduction.— Starting with an initial state and the microscopic Hamiltonian of an isolated system governing its dynamics, it might seem that the unitary evolution of the system would prevent thermalization. This puzzle is addressed by the Eigenstate Thermalization Hypothesis (ETH) [1, 2, 3, 4, 5]. According to the ETH, for non-integrable systems, the high-energy states and the observables of interest will eventually converge to a description in accordance with equilibrium statistical mechanics, achieving thermal equilibrium. In contrast to the ETH, many-body localized systems defy this trend due to their emergent integrability [6]. Systems exhibiting Quantum Many-Body Scars (QMBS) offer a contrasting example [7, 8, 9]. They also evade ETH, but only in an exponentially small fraction of states in the Hilbert space.
@@
The effect of QMBS was first observed in a Rydberg-atom quantum simulator, marked by persistent revivals for particular initial states, in contrast to the vast majority of other high-energy initial states [10, 11]. In parallel with the exact construction of highly excited eigenstates in a non-integrable model [12, 13], the experiment has sparked intense research across different systems, where a weak breaking of the ETH takes place [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. In this case, the presence of few anomalous eigenstates, can leave an imprint on thermalization. Among their special features is that they are characterized by atypically low entanglement entropy, compared with other states arbitrarily close in the spectrum. This is in stark contrast to what is expected from the ETH. In particular, QMBS may exhibit an area law. This means that the entanglement between a subsystem and its complement is proportional to the area of the boundary that divides them. In turn, highly excited states are expected to have an entanglement that grows linearly with the volume of the subsystem.
@@
The interplay between gauge symmetry and scarring phenomena is not fully understood. Gauge theories naturally lead to constraints, and specific constraints can give rise to QMBS, as observed in the PXP model [10, 27, 28]. Gauge theories also feature local symmetries, which segment the Hamiltonian into distinct sectors that do not mix under time evolution. Supplied by a condition on physical states, this leads to local constraints in the form of Gauss’ law. Beyond the PXP model, QMBS have also been observed in other gauge theories, such as the Abelian case in the presence of matter in $1+1$D [29], pure gauge in $2+1$D [30, 31, 32] and non-Abelian in $2+1$D [33]. So far, the research on QMBS in Abelian pure gauge theories has exclusively focused on spin-$1/2$ Quantum Link Models (QLM) [30, 31, 32], which also exhibit other interesting phenomena such as crystalline confining phases or confining strings with fractionalized electric flux stands [34, 35, 36, 37]. This Letter demonstrates, for the first time, the existence of QMBS in link models with arbitrary integer spins.
@@
Addressing questions related to real-time dynamics falls outside the capabilities of conventional lattice gauge theory methods, as Monte Carlo simulations face inefficiencies due to severe sign problems associated with real-time processes (see e.g. [38, 39]). Sign problems have motivated the development of the field of quantum simulations of gauge theories [40, 41, 42, 43, 44, 45]. Formulations like QLMs are essential for preserving gauge symmetry while ensuring a finite Hilbert space for a finite lattice volume. Recovering the theory in the continuum limit may take different routes [46, 47, 48, 49, 50, 51, 52]. No matter the path, the road ahead is long until we can make decisive contributions to theories like Quantum Chromodynamics (QCD). Hence, it is imperative to pinpoint intriguing phenomena that can drive experiments along this path. The study of QMBS is a meritorious example as it is present in simple gauge theories and probes fundamental aspects of quantum many-body theory.
@@
Demonstrating the presence of QMBS for large Hilbert spaces per gauge link, beyond spin-$1/2$, serves several fundamental purposes: it directly addresses how widespread QMBS are across many-body systems; it reveals novel mechanisms for their formation, illuminating the role played by gauge symmetry; and it helps to guide quantum simulation experiments toward interesting questions before the necessary complexity of QCD is achieved.
@@
In this Letter, we demonstrate the extensive presence of QMBS across $2+1$ dimensional $U(1)$ gauge theories without matter. Concretely, we explicitly construct mid-spectrum states that satisfy area law entanglement for arbitrary integer spin, specifically for a simply truncated Hilbert space per link. We further verify the existence of these states by numerically determining the system’s eigenstates and calculating their entanglement and Shannon entropy for spin 1 and 2. Our findings unveil the presence of QMBS for single-leg ladders of the spin-1 QLM, which are more accessible to experiments than larger systems. Furthermore, our analytical approach allows us to identify QMBS for arbitrary integer spin and system size, circumventing the limitations of existing numerical methods when applied to large dimensional Hilbert spaces.
@@
Link models in $2+1$D.— We consider models on square $L_{1}\times L_{2}$ lattices with bosonic degrees of freedom living on the links
@@
where the indices $n\equiv\left(n_{1},n_{2}\right)$ represent lattice sites, while the labels $i\in\left\{1,2\right\}$ represent the two directions. The first terms of the Hamiltonian are plaquette terms constructed by acting on each of the four links of a plaquette with operators represented by $U$. The term $V$, which we will call generically potential, will always be diagonal in the electric field $E_{ni}$. The commutation relations between these variables are $\left[E_{mi},U_{nj}\right]=U_{mi}\delta_{mn}\delta_{ij}$. Under these general conditions, the Hamiltonian has a set of local symmetries. Concretely, there is one generator of local gauge transformations $G_{n}$ per lattice site, which commutes with the Hamiltonian
@@
In Truncated Link Models (TLM) [29] the electric field is simply truncated, meaning $U\left|S\right>=0$, $U^{\dagger}\left|-S\right>=0$ and $U\left|n\right>=\ket{n+1}$ otherwise. $U_{ni}$ and $U^{\dagger}_{ni}$ still do not commute, though the non-zero commutation relation is removed to the edge of the local Hilbert space (to the states |$\ket{-S}$ and |$\ket{S}$). With this construction, there will be $12S+$ 1 states per link. We generically refer to $S$ as the total spin, irrespective of the formulation we are using. The $S=1$ TLM is equivalent to the $S=1$ QLM.
For the Hamiltonian (1) we will also write
@@
where $H^{-}_{n}=U_{n1}^{\dagger}U_{n+\hat{1}2}^{\dagger}U_{n2}U_{n+\hat{2}1}$ and $H^{+}=\left(H^{-}\right)^{\dagger}$ are kinetic operators. This notation follows the convention that we will adopt throughout the letter. “Plaquette at site $n$” refers to the plaquette for which $n$ is its lower left vertex. In order to simplify the formulas we will also adopt the convention $\left(H^{+}\right)^{-m}\equiv\left(H^{-}\right)^{m}$. This should not be confused with the inverse of $\left(H^{+}\right)^{m}$, since these operators are not invertible.
Zero-mode scars.— We refer to zero-modes as eigenstates of the plaquette term of the Hamiltonian with zero eigenvalues. In our notation, these states satisfy
@@
Zero-modes can play a crucial role in forming quantum many-body scars. Systems with a spectral symmetry, together with point-group symmetries, can exhibit an exponential number of zero-modes [56]. Their existence follows from an index theorem, which we review and adapt to our models of interest in the Supplementary Material [53]. While one might expect these eigenstates to be thermal, it has been shown that low entangled states can be constructed within this subspace for specific cases. It was further conjectured that this is a generic property of local Hamiltonians with an exponential number of zero-modes [57]. Such a result was also observed in the case of gauge theories in the spin-$1/2$ QLM, even if it is possible to find scars that are not constructed exclusively from zero-modes [31, 32].
We show that it is also possible to construct such states for arbitrary spin truncated link models. The number of these low entropy states grows exponentially with volume. By choosing appropriately the potential term of the Hamiltonian, different linear combinations can be isolated as low entropy eigenstates. To establish these states as scars, one should show that the model in question is non-integrable. It is expected that this is the case for pure gauge theories in (2+1)-d, and we further demonstrate this for a single-leg ladder with $S=1$ in the Supplementary Material [53].
QMBS in TLM with arbitrary integer spin.— Next, we show that zero-modes with area-law entanglement entropy exist in truncated link models of all integer spins. We start by partitioning the 2D lattice into $2\times 1$ and $1\times 2$ tiles, as depicted on the left of Fig. 1. A given partition will be called tiling and will be denoted by $T$. It can be represented by a set of tuples $\left(n,n^{\prime}\right)$ where the two entries indicate the two plaquettes making up each tile. We then define the states
@@
where |$\ket{\mathbf{0}}$ is the state where all links have value zero, and $0\leq i\leq S$. The states |$\ket{\psi_{s}^{(i,T)}}$ have the property $(H^{-}_{m}+H^{+}_{m^{\prime}})\ket{\psi_{s}^{(i,T)}}=(H^{+}_{m}+H^{-}_{m^{\prime}})\ket{\psi_{s}^{(i,T)}}=0$ if $(m,m^{\prime})$ is a tile in $T$. Therefore,
@@
making them zero-modes.
Figure 1: Left: Example of a tiling $T$ of the 2D lattice, consisting of partitioning it into $2\times 1$ and $1\times 2$ tiles, to a total of 8 tiles. Right: Representation of a scar state based on an alternative tiling and in the dual representation. An independent sum over every $k_{i}$ with a weight of $\prod_{i}\left(-1\right)^{k_{i}}$ is assumed. See Eq. (Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories).
@@
In the dual representation, the state (Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories) in a $2\times 1$ plaquette system is
@@
All states that contribute in (Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories) have a height variable representation where there is one state out of the sum (7) on all tiles of the tiling $T$. The state sums over all $(S+1)^{\left|T\right|}$ combinations of these states. A pictorial representation of this tiling is shown on the right-hand side of Fig. 1. This tiling structure is similar to the lego and sub-lattice scars constructed for the $S=1/2$ case [31, 32], and allows for the construction of an exponential number of scars for arbitrary truncated link models. A formal definition of this tiling, in the form of a tiling product, can be found in the Supplementary Material [53]. There, we also describe a more general framework to construct more low-entropy zero-modes in TLMs.

### v3

@@ Abstract
The existence of Quantum Many-Body Scars, which prevents thermalization from certain initial states after a long time, has been established across different quantum many-body systems. These include gauge theories corresponding to spin-1/2 quantum link models. Establishing quantum scars in gauge theories with high spin is not accessible with existing numerical methods, which rely on exact diagonalization. We systematically identify scars for pure gauge theories with arbitrarily large integer spin $S$ in $2+1$D, where the electric field is restricted to $12S+$ 1 states per link. Through an explicit analytic construction, we show that the presence of scars is widespread in $2+1$D gauge theories for arbitrary integer spin. We confirm these findings numerically for small truncated spin and $S=1$ quantum link models. Our analytic construction establishes the presence of scars far beyond volumes and spins that can be probed with existing numerical methods and can guide quantum simulation experiments toward interesting non-equilibrium phenomena, inaccessible otherwise.
I Introduction
@@
Evolving an isolated quantum system over long times might suggest that the unitary evolution of the system would prevent thermalization. This puzzle is addressed by the Eigenstate Thermalization Hypothesis (ETH) for quantum many-body systems [1, 2, 3, 4, 5]. According to the ETH, for non-integrable systems, the high-energy states and the observables of interest will eventually converge to a description following equilibrium statistical mechanics, achieving thermal equilibrium. In contrast to the ETH, many-body localized systems defy this trend due to their emergent integrability [6]. Systems exhibiting Quantum Many-Body Scars (QMBS) offer a contrasting example [7, 8, 9]; they also evade the ETH, but only in an exponentially small fraction of states in the Hilbert space.
@@
The effect of QMBS was first observed in a Rydberg-atom quantum simulator, marked by persistent revivals for particular initial states, in contrast to the vast majority of other high-energy initial states [10, 11]. In parallel with the exact construction of highly excited eigenstates in a non-integrable model [12, 13], the experiment has sparked intense research in a variety of quantum many-body systems, where a weak breaking of the ETH takes place [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. In these systems, the presence of few anomalous eigenstates can leave an imprint on thermalization. Among the special features of the anomalous states is that they are characterized by atypically low entanglement entropy, compared with other states arbitrarily close in the spectrum. In particular, QMBS may exhibit an area law of bipartite entanglement entropy. This means that the entanglement between a subsystem and its complement is proportional to the area of the boundary that divides them. This contradicts the ETH, which predicts that highly excited states have an entanglement that grows linearly with the volume of the subsystem.
@@
The interplay between gauge symmetry and scarring phenomena is not fully understood. Gauge theories naturally lead to constraints, and specific constraints can give rise to QMBS, as observed in the PXP model [10, 27, 28]. The constraints in gauge theories stem from local symmetries, which segment the Hamiltonian into distinct sectors that do not mix under time evolution. Supplied by a condition on physical states, this leads to local constraints in the form of Gauss’ law. Beyond the PXP model, QMBS have also been observed in other gauge theories, such as the Abelian case in the presence of matter in $1+1$D [29], pure gauge in $2+1$D [30, 31, 32] and non-Abelian in $1+1$D [33] and $2+1$D [34]. So far, the research on QMBS in Abelian pure gauge theories has exclusively focused on spin-$1/2$ Quantum Link Models (QLM) [30, 31, 32], which also exhibit other interesting phenomena such as crystalline confining phases or confining strings with fractionalized electric flux stands [35, 36, 37, 38]. This work demonstrates, for the first time, the existence of QMBS in link models with arbitrary integer spins.
@@
Addressing questions related to real-time dynamics falls outside the capabilities of conventional lattice gauge theory methods, as Monte Carlo simulations face inefficiencies due to severe sign problems (see e.g. [39, 40]). Sign problems have been a driving force in the development of quantum simulations of gauge theories [41, 42, 43, 44, 45, 46, 47, 48]. Formulations like QLMs are essential for preserving gauge symmetry while ensuring a finite Hilbert space for a finite lattice volume. Recovering the original theory may take different routes [49, 50, 51, 52, 53, 54, 55]. Regardless of the chosen method, significant contributions to theories such as Quantum Chromodynamics (QCD) are still far off. Therefore, it is crucial to identify intriguing phenomena that can guide experimental efforts in this direction. The study of QMBS is a meritorious example, as they appear in simple gauge theories and probe fundamental aspects of quantum many-body theory.
@@
Demonstrating the presence of QMBS for large Hilbert spaces per gauge link, beyond spin-$1/2$, serves several fundamental purposes: it directly addresses how widespread QMBS are across many-body systems; it reveals novel mechanisms for their formation, illuminating the role played by gauge symmetry; and it helps to guide quantum simulation experiments toward interesting questions before the complexity of QCD is achieved.
@@
In this work, we demonstrate the extensive presence of QMBS across $2+1$D $U(1)$ gauge theories without matter. Concretely, we explicitly construct mid-spectrum states that satisfy area law entanglement for arbitrary integer spin, specifically for a simply truncated Hilbert space per link. We further verify the existence of these states by numerically determining the system’s eigenstates and calculating their entanglement and Shannon entropy for spin 1 and 2. Our findings unveil the presence of QMBS for single-leg ladders of the spin-1 QLM, which are more accessible to experiments than wider systems. Furthermore, our analytical approach allows us to identify QMBS for arbitrary integer spin and system size, circumventing the limitations of existing numerical methods when applied to large dimensional Hilbert spaces.
II Link models in $2+1$D
@@
We consider models on square $L_{1}\times L_{2}$ lattices with bosonic degrees of freedom living on the links
@@
where the indices $n\equiv\left(n_{1},n_{2}\right)$ represent lattice sites, while the labels $i\in\left\{1,2\right\}$ represent the two directions. The first terms of the Hamiltonian are plaquette terms, constructed by acting on each of the four links of a plaquette with operators represented by $U$. The term $V$, which we will call generically potential, will always be diagonal in the electric field $E_{ni}$. The commutation relations between these variables are $\left[E_{mi},U_{nj}\right]=U_{mi}\delta_{mn}\delta_{ij}$. Under these general conditions, the Hamiltonian has a set of local symmetries. Concretely, there is one generator of local gauge transformations $G_{n}$ per lattice site, which commutes with the Hamiltonian
@@
In Truncated Link Models (TLM) [59] (see also [29, 60]) the electric field is simply truncated, meaning $U\left|S\right>=0$, $U^{\dagger}\left|-S\right>=0$ and $U\left|n\right>=\ket{n+1}$ otherwise. $U_{ni}$ and $U^{\dagger}_{ni}$ still do not commute, though the non-zero commutation relation is removed to the edge of the local Hilbert space (to the states |$\ket{-S}$ and |$\ket{S}$). With this construction, there will be $12S+$ 1 states per link. We generically refer to $S$ as the total spin, irrespective of the formulation we are using. The $S=1$ TLM is equivalent to the $S=1$ QLM, up to normalization of Hamiltonian parameters.
We will also write the Hamiltonian (1) as
@@
where $H^{-}_{n}=U_{n1}^{\dagger}U_{n+\hat{1}2}^{\dagger}U_{n2}U_{n+\hat{2}1}$ and $H^{+}=\left(H^{-}\right)^{\dagger}$ are kinetic operators. This notation follows the convention that we will adopt throughout the paper. “Plaquette at site $n$” refers to the plaquette for which $n$ is its lower left vertex. In order to simplify the formulas we will also adopt the convention $\left(H^{+}\right)^{-m}\equiv\left(H^{-}\right)^{m}$. This should not be confused with the inverse of $\left(H^{+}\right)^{m}$, since these operators are not invertible.
To demonstrate the existence of scars in these models, one should show that they are non-integrable. It is expected that this is the case for pure gauge theories in $2+1$D, and we will further demonstrate it here for a single-leg ladder with $S=1$.
II.1 Non-Integrability of the Spin-1 Ladder
We consider a single row of plaquettes with spin $S=1$ links, open boundary conditions in both directions and the potential $V=\lambda\sum_{n_{1}=0}E_{\left(n_{1},1\right)1}$. We show that his model is non-integrable by studying the statistics of the spectrum between adjacent energy levels. We follow the method introduced in [61], which was also applied to QLMs in [30, 31]. This will be one of the many cases in which we demonstrate the existence of mid-spectrum low entropy states, establishing the presence of quantum many-body scars. While we do not demonstrate the non-integrability of all models considered, the result is expected to extend beyond this case.
We start by resolving the symmetries of the system. Translation symmetry is broken due to the open boundaries. The potential breaks charge conjugation symmetry. Reflection across the horizontal axis, as defined above, is identical to charge conjugation and is therefore also broken by the potential. Reflection symmetry with respect to the vertical axis needs to be resolved. This results in two sectors, labeled with the eigenvalue ±$\pm 1$. The distribution $p\left(r\right)$ of consecutive level spacing ratios
$r_{n}=\min\left\{\frac{E_{n+1}-E_{n}}{E_{n}-E_{n-1}},\frac{E_{n}-E_{n-1}}{E_{n+1}-E_{n}}\right\}$ (4)
of both sectors match the Gaussian Orthogonal Ensemble (GOE)
$p_{GOE}(r)=\frac{27}{4}\frac{r+r^{2}}{(1+r+r^{2})^{5/2}}$ (5)
as seen in Fig. 1. This is in contrast to the Poisson distribution expected for integrable models
$p_{P}(r)=\frac{2}{(1+r)^{2}},$ (6)
which is also plotted for reference.
Figure 1: Level spacing distribution for a $12\times 1$ ladder with the height potential at varying $\lambda$. Left: +1 reflection symmetry sector. Right: -1 reflection symmetry sector. Both match the GOE distribution well, indicating that this system is not integrable.
III Zero-mode scars
We refer to zero-modes as eigenstates of the plaquette term of the Hamiltonian with zero eigenvalues. In our notation, these states satisfy
@@ III Zero-mode scars
Zero-modes can play a crucial role in forming quantum many-body scars. Systems with a spectral symmetry, together with point-group symmetries, can exhibit an exponential number of zero-modes [62]. Their existence follows from an index theorem, which we review and adapt to our models of interest in the Supplementary Material [56]. While one might expect these eigenstates to be thermal, it has been shown that states with low entanglement can be constructed within this subspace for specific cases. It was further conjectured that this is a generic property of local Hamiltonians with an exponential number of zero-modes [63]. This property was observed explicitly in the case of gauge theories in the spin-$1/2$ QLM, even if it is possible to find scars that are not constructed exclusively from zero-modes [31, 32].
We show that it is also possible to construct such states for arbitrary spin truncated link models. The number of these low entropy states grows exponentially with the volume. By choosing the potential term of the Hamiltonian appropriately, different linear combinations can be isolated as low entropy eigenstates.
III.1 QMBS in TLM With Arbitrary Integer Spin
Next, we show that zero-modes with area-law entanglement entropy exist in truncated link models of all integer spins. We start by partitioning the 2D lattice into $2\times 1$ and $1\times 2$ tiles, as depicted on the left of Fig. 2. A given partition will be called tiling and will be denoted by $T$. It can be represented by a set of tuples $\left(n,n^{\prime}\right)$ where the two entries indicate the two plaquettes making up each tile. We then define the states
@@ III.1 QMBS in TLM With Arbitrary Integer Spin
where |$\ket{\mathbf{0}}$ is the state where all links have value zero, and $0\leq i\leq S$. The states |$\ket{\psi_{s}^{(i,T)}}$ have the property
$(H^{-}_{m}+H^{+}_{m^{\prime}})\ket{\psi_{s}^{(i,T)}}=(H^{+}_{m}+H^{-}_{m^{\prime}})\ket{\psi_{s}^{(i,T)}}=0$ (9)
if $(m,m^{\prime})$ is a tile in $T$. Therefore,
@@
making them zero-modes for truncated link models of arbitrary spin. For quantum link models with spin $S>1$, this would no longer be the case due to the electric field-dependent pre-factors when raising/lowering the spins.
Figure 2: Left: Example of a tiling $T$ of the 2D lattice, consisting of partitioning it into $2\times 1$ and $1\times 2$ tiles, to a total of 8 tiles. Right: Representation of a scar state based on an alternative tiling and in the dual representation. An independent sum over every $k_{i}$ with a weight of $\prod_{i}\left(-1\right)^{k_{i}}$ is assumed. See Eq. (24).
@@
In the dual representation, the state (24) in a $2\times 1$ plaquette system is
@@
All states that contribute in (24) have a height variable representation where there is one state out of the sum (11) on all tiles of the tiling $T$. The state sums over all $(S+1)^{\left|T\right|}$ combinations of these states. A pictorial representation of this tiling is shown on the right-hand side of Fig. 2. This tiling structure is similar to the lego and sub-lattice scars constructed for the $S=1/2$ case [31, 32], and allows for the construction of an exponential number of scars for arbitrary truncated link models. A formal definition of this tiling, in the form of a tiling product, can be found in the Supplementary Material [56]. There, we also describe a more general framework to construct more low-entropy zero-modes in TLMs.
III.2 Beyond Zero-Mode Building Blocks
The above-described mechanism follows two basic ingredients. It constructs zero-modes from a basic tile made of two plaquettes joined together in one link. The full state is obtained by tiling these building blocks together. In this section, we demonstrate that it is possible to form zero-modes with different structures. This demonstrates that even though we have constructed numerous scars for arbitrary spin, there are additional routes to obtain other types of zero-mode scars. We provide explicit constructions for the $S=1$ case. The two examples described here exhibit a sublattice structure similar to what has been found for spin-$1/2$ [31, 32]. One can be constructed by making use of diagonal tiles, while the other does not exhibit any tiling structure. A third example, that we have found for the specific case of a $4\times 4$ volume, can be found in the Supplemental Material [56]. To the best of our knowledge, that specific construction is not generalizable for arbitrary volumes.
Diagonal Tiling
We consider a scar with the generic form previously described in (24) for the specific case of spin $S=1$, i.e.
[…]
in the dual representation. The two plaquettes that make up tiles no longer share a link, but neighbor each other diagonally in a pattern depicted in Fig. 4. Individual tiles are not zero-modes, i.e. they do not satisfy (9). However, the resulting state is a zero-mode of the full system. This follows because each plaquette neighbors both sites of another tile. One of the plaquettes of this tile will take the value ±$\pm 1$. None of the plaquettes can therefore ever take height $\mp 1$.
Non-Tiling Scar
It is also possible to have scars that do not follow (24) nor exhibit a similar tiling structure. We divide the lattice into two sublattices $A$ and $B$ and construct a zero-mode according to
|$\ket{\psi_{s}}=\frac{1}{2^{L_{1}L_{2}/2}}$ $\prod_{m\in A}\left(1-\left(H^{+}_{m}\right)^{2}P^{+}_{m}-\left(H^{-}_{m}\right)^{2}P^{-}_{m}\right)$
$\prod_{n\in B}\left(H^{+}_{n}-H^{-}_{n}\right)\ket{\mathbb{0}},$ (13)
where $P^{\pm}_{m}$ is the projector on ±$\pm 1$ on all plaquettes surrounding the plaquette at $m$. An illustration can be found in Fig. 4.
Refer to caption         Refer to caption
Figure 4: Illustrations of $S=1$ zero-mode scars that are eigenstates of $\sum_{n}E_{n}^{2}$, exhibiting sublattice structure. Left: Scar built by singlet states of two plaquettes connected diagonally. Plaquettes connected by red or green lines form tiles. Right: Singlet states are placed in sublattice $B$ while sublattice $A$ is filled by 0’s. If all plaquettes surrounding a 0 have the same value of ±$\pm 1$, then that plaquette is combined with ±$\pm 2$.
This state can be understood by considering one sublattice and creating a linear superposition of heights ±$\pm 1$. In our example, this is done for sublattice $B$. For plaquettes of sublattice $A$, the height variable is stuck at zero as long as all surrounding plaquettes are not equal. Otherwise, the plaquette can either be raised or lowered depending on the value of the surrounding plaquettes being positive or negative. By combining those states with the respective lowered and raised states with the opposite sign, we obtain a zero-mode.
IV Numerical results and discussion